# Randomness in atomic disorder and consequent squandering of spin-polarization in a ferromagnetically fragile quaternary Heusler alloy FeRuCrSi

### v1

@@ Abstract
Ru2-xFexCrSi ( 0 $<$ x $<$1) system is theoretically predicted to be one of the very few known examples of robust half-metallic ferromagnet with 100% spin polarization. Since Cr is considered to be the main contributor to magnetism, the Fe/Ru substitution is not expected to disturb its magnetic properties any significantly, and hence all Fe-containing members of the series are predicted to follow Slater-Pauling rule with a saturation magnetic moment of 2 ${\mu_{B}}$/f.u. However, contrarily to the theoretical expectations, some experiments rather show a linear variation of the saturation magnetization and Curie temperature with Fe (x) substitution. The equiatomic member FeRuCrSi of this family is also considered as a technologically important material, where the band structure calculations suggest the material to be spin gapless semiconductor. Through our in-depth structural analysis of FeRuCrSi using X-ray diffraction, extended X-ray absorption fine structure and 57Fe Mössbauer spectrometry, we found a random disorder between Fe and Ru sites, while the magnetic moment in this system is actually contributed by Fe atoms, questioning the very basic foundation of the half-metallic character proposed by all theoretical calculations on Ru2-xFexCrSi series. Our Mössbauer result also envisions a rather rare scenario where the main physical properties are intricately correlated to the chemistry of the material in the form of random atomic disorder on a localised scale.
@@ I Introduction
Heusler alloys have recently garnered attention due to their diverse properties, including half-metallicity [10, 11, 12, 13, 14], spin gapless semiconductivity [15, 16, 17], spin-semimetals [18, 19], spin glass [20], re-entrant spin glass [21], bipolar magnetic semiconducting characteristics [22], etc. The exploration of new Heusler alloys is crucial for advancing this area of study, as despite theoretical predictions [23, 24, 25, 26] of numerous Heusler alloys possessing SGS features, only a limited number of experimental realizations have been achieved so far [27, 15, 16, 17]. Among them, Mn2CoAl [27], CoFeMnSi [15], CoFeCrGa [16], and CoFeMnSn [17] have been experimentally established as SGS materials, with the additional discovery of fully compensated ferrimagnetic SGS in CrVTiAl [28]. It is pertinent to note here that the most of the Heusler alloys often form with inherent and irrepressible structural disorder which is generally argued to be responsible for the nonconformity with the theoretically predicted HMF and SGS properties. So, finding compound having robust HMF and SGS properties will be an effective solution for this kind of situation.
@@
Ru2-xFexCrSi is one such system where first principles band structure calculations predict robust HMF properties against structural disorder throughout the series [29, 30, 31]. The equiatomic member of the series, FeRuCrSi, has been additionally predicted to exhibit SGS characteristics [32, 33, 34]. The quaternary compound FeRuCrSi can be thought of as the intermediate between two ternary Heusler alloys Fe2CrSi [35] and Ru2CrSi [36]. Since Ru and Fe are isoelectronic atoms, the total valence electron count (VEC) of all members of Ru2-xFexCrSi remain invariant (VEC = 26) irrespective to the value of x. Following the Slater-Pauling (S-P) rule [37, 38] for HMF Heusler alloy, the total magnetic moment for any member of the series is expected to be M = (VEC-24) = 2 $\mu_{B}$/f.u. However, the Ru-rich end-composition, Ru2CrSi is experimentally found to undergo antiferromagnetic transition below TNsimilar-$\sim$14 K [36]. Subsequent theoretical calculation indeed shows lower ground state energy for the antiferromagnetic spin arrangement in this compound [39]. As Fe concentration is gradually increased, the system initially transforms to a spin-glass state, before it finally settles to a ferromagnetic ground state for x $>$ 0.3 [40, 41, 42, 31, 43]. Fe2CrSi, the other-end member of the series, exhibit Curie temperature at 520 K with saturation magnetic moment of 2 $\mu_{B}$/f.u., in line with the S-P rule [35]. Theoretical calculations suggests Cr-atoms are the major contributor of magnetism in this series of compounds, although this postulation is yet to be experimentally verified for compositions with x $\neq$ 0. Interestingly, some of the experimental reports rather suggest that a linear variation of magnetic moment occur with Fe substitution in Ru2-xFexCrSi, contrary to the invariance of magnetic moment expected from the S-P rule for HMF systems [40]. Additionally, even for FeRuCrSi, large variations in its basic properties like Curie temperature (TC) and saturation magnetization (Msat) at low temperature have been reported in the literature [40, 33]. Thus, the members of Ru2-xFexCrSi series not only sometimes fail to conform to the theoretical prediction, the experimental results on TC and MSat reported by different groups on the same compositions are also not reproducible. As mentioned earlier, since the Heusler alloys are known to form with inherent structural disorder, a priori one would expect that such variations in properties reported by various groups are related to different natures and extents of disorder present in the respective materials, despite having the same nominal composition. In this work, we have focussed on the equiatomic FeRuCrSi, and investigated the nature of disorders to explain the above-mentioned discrepancies among different experimental results as well as between experiment and theory. Very surprisingly, our investigation reveals that completely contrary to the theoretically established magnetic moment on Cr in FeRuCrSi, the magnetisam is actually contributed by Fe atoms. The large variation in TC and Msat can also be explained by the local arrangement of Fe/Ru in each unit cells. Our work thus firmly establishes not only the role of anti-site disorders, but the effect of randomness of Fe/Ru distribution in this compound as well.
@@ II.1 Experimental
Extended X-ray absorption fine structure (EXAFS) measurements were conducted at the Energy Scanning EXAFS beamline (BL-9) at the Raja Ramanna Centre for Advanced Technology (RRCAT) in Indore, India, employing standard EXAFS measurement protocols [13, 21]. Magnetic measurements were carried out in the temperature range of 3–380 K under magnetic fields up to 70 kOe using a commercial SQUID-VSM instrument (Quantum Design Inc., USA). 57Fe transmission Mössbauer spectrometry was conducted at both room temperature (300 K) and liquid nitrogen temperature (77 K) to investigate the local atomic environments of Fe atoms in the sample. The samples consisted of a thin powder layer containing about 5 mg Fe/cm2. Spectra were collected using an electromagnetic transducer with a triangular velocity profile and a 57Co source diffused into a Rh matrix within a bath cryostat. The hyperfine structures were modeled using a least-square fitting procedure, involving quadrupolar doublets composed of Lorentzian lines, utilizing the in-house program ‘MOSFIT’. Isomer shift values were reported with reference to ${\alpha}$–Fe at 300 K, and the velocity was controlled using a standard of ${\alpha}$–Fe foil.
@@ II.2 Computational
The enthalpy of formation, electronic structure, and spin polarization at ground state were determined using Density Functional Theory (DFT) calculations with the projector augmented wave (PAW) method [45], as implemented in the Vienna ab initio simulation package (VASP) [46, 47]. The exchange-correlation was modeled by the generalized gradient approximation modified by Perdew, Burke, and Ernzerhof (GGA-PBE) [48]. Detailed information on calculation parameters can be found in previous literatures [12, 13, 14, 21].
@@ III.1 Structure optimization and electronic structure calculations
Table 1: Calculated enthalpy of formation (${\Delta_{f}{H}}$) for 3 different ordered atomic arrangement of FeRuCrSi.
4a 4b 4c 4d $\Delta_{f}{H}$ $(kJ/molbyatoms)$
Type-1 Si Cr Ru Fe -33.63
Type-2 Si Ru Cr Fe -4.04
Type-3 Si Fe Cr Ru -29.25
A quaternary Heusler alloy, denoted as ${XX^{\prime}YZ}$, possesses four distinct crystallographic sites: 4a (0,0,0), 4b (0.5,0.5,0.5), 4c (0.25,0.25,0.25), and 4d (0.75,0.75,0.75). Generally, the sp-element Z occupies the 4a position, while the transition metals reside in the remaining three sites. Although there are six possible configurations, permutations of atoms in the 4c and 4d positions result in energetically equivalent configurations. Consequently, only three distinct configurations (Type-1, 2, and 3) were considered, as presented in Fig. 1 and Table 1. The relative stabilities of these configurations were assessed through DFT calculations by comparing their enthalpies of formation.
@@
Consistent to that expected for HMF systems with VEC 26 [37, 49, 38], our calculation also yields the magnetic moment to be 1.943 ${\mu_{B}}$. Moreover, atom-specific magnetic moments were computed, revealing values of Fe = 0.146 $\mu_{B}$, Ru = -0.18 $\mu_{B}$, Cr = 1.984 $\mu_{B}$, and Si = -0.007 $\mu_{B}$ for the ordered structure, conforming with earlier published reports [34, 32, 33]. The theoretical analysis suggests that Cr significantly contributes to the magnetic moment, and the moment on the Fe site rather exhibits an induced character. Similar behaviour has earlier been observed in FeMnVAl  [12] and FeMnVGa  [13] where magnetic atom Mn and nonmagnetic Fe are distributed randomly in the 4c and 4d sites, although in FeRuCrSi, Cr occupies 4b sites whereas Ru and Fe occupy 4c and 4d positions, respectively.
@@ III.2 X-ray diffraction
Figure 3: Rietveld refinement of the powder XRD pattern of FeRuCrSi (disordered structure) measured at room temperature. Inset (i) shows the magnified view of the XRD data for as-cast and annealed compound, where the impurity peaks in the annealed sample are marked by asterisk sign. Inset (ii) and (iii) show the magnified view of the Rietveld refinement results considering the ordered structure and disordered structure, respectively.
@@
We have carried out XRD measurements at room temperature for both as-cast as well as annealed sample. We found a few additional small peaks in the XRD patterns of the annealed sample which are absent in the XRD pattern of the as cast samples and also not conforming to the cubic character of Heusler alloy (inset (i) of Fig. 3). The result thus suggest that annealing tend to introduce secondary phase in the system. Consequently, all further measurements are carried out on as-cast compound only. Fig. 3 represent the Rietveld refinement of the XRD pattern of as-cast FeRuCrSi. Theoretical calculations suggests that ordered Type-1 structure possess minimum energy compared to other ordered structures. However, the earlier published experimental results on the crystal structure of FeRuCrSi are quite contradictory. Although some of the reports suggested the ordered LiMgPdSn structure-type [40], a few others also reported disordered atomic arrangements without determining the exact nature of disorder [33].
@@
Our attempt to fit the XRD pattern using the ordered Type-1 structure model however failed to properly address the superlattice (111) Bragg peak intensity (inset (ii) of Fig. 3). For Heusler systems, the (111) and (200) Bragg peak intensities are considered to be indicative of structural disorder present in the system. For example, in A2-type (random mixing of elements in 4a, 4b, 4c and 4d sites) disorder, both the peaks found to be absent whereas in B2-type (cross-substitution between 4a & 4b and between 4c & 4d) only the (111) peak vanishes. In our case, both the 111 and 200 lines are present, thus ruling out A2 and B2-type disorder (inset (iii) of Fig. 3). Nonetheless, the intensity of 111 line is substantially reduced when compared to the theoretically estimated one for an ordered Type-1 structure. The reduction in intensity of (111) peak, however, could be accounted for if the disorder is limited to the 4c and 4d sites only, where Fe and Ru atoms are distributed in 50:50 ratio, averaging over the sample volume. This type of disorder earlier observed in CoRuMnSi [50], FeMnVAl [12], FeMnVGa [13] and NiRuMnSn [21]. Since the x-ray scattering cross-section of Cr and Fe are very similar, XRD remains insensitive to any further disorder between Cr (4b) and Fe (4c/4d) sites. The lattice parameter is found to be 5.757  Å, and matches with most of the other reported values [40, 34].
@@ III.3 EXAFS
The simultaneous fitting of Fe and Ru edges could only be achieved under the assumption of disorder, confirming disorder in the examined compound based on the XRD data. The best-fit theoretical spectra and experimental data are presented in Fig. 5, and the corresponding parameters are listed in Table 2. The EXAFS estimate a ratio of Fe/Ru mixing as 59:41 and 45:55, while measuring from Fe and Ru edges, respectively. The EXAFS data fitting aligns well with a disordered structure of Type-1, corroborating our XRD data analysis.
III.4 Electronic structure calculations for disordered crystal structure
Figure 6: Partial density of states for FeRuCrSi for disordered structure.
Following our finding of a 50:50 site disorder between Fe and Ru atoms, we have extended the electronic structure calculations for FeRuCrSi considering the same disorder in the theoretical analysis (Fig. 6). Theoretical computations reveal that despite structural disorder, the system maintains 100% spin polarization. Additionally, the total magnetic moment estimated to be 1.94 $\mu_{B}$, consistent with both the ordered structure and the Slater-Pauling value. Specifically, magnetic moments for individual sites in the disordered structure are estimated as follows: Si = -0.008 $\mu_{B}$, Cr = 1.898 $\mu_{B}$, Ru = -0.161 $\mu_{B}$, and Fe = 0.211 $\mu_{B}$.
III.5 Magnetic properties
Figure 7: Temperature dependence of magnetization in FeRuCrSi measured in 100 Oe applied magnetic field under ZFC (blue circle) and FC (red circle) conditions. Curie temperature, TC, is determined from the minima in dM/dT vs. T plot. Inset shows inverse susceptibility data measured under FC condition. At high temperature region, Curie-Weiss fit with extrapolated section has also been shown.
Fig. 7 represents the magnetization vs. temperature of FeRuCrSi measured in both zero field cooled (ZFC) and field cooled (FC) condition under an application of 100 Oe field. Thermo-magnetic irreversibility observed between ZFC and FC magnetization which vanishes when the external field strength is enhanced to 500 Oe and beyond. The susceptibility shows a sharp ferromagnetic to paramagnetic transition near its Curie temperature, TC = 217 K, which is determined from the minima of d(M/H)/dT curve.

### v2

@@ Abstract
The FexRu2-xCrSi ($20<x<$ 2) system should theoretically reckoned to be one of the few known examples of a robust half-metallic ferromagnet with 100% spin polarisation, with Cr atoms deemed to be the main contributor to magnetism. In this system, the FeRuCrSi equiatomic member is considered a technologically important material, since band structure calculations suggest that it is a spin-gap-free semiconductor. Through our extensive structural analysis of FeRuCrSi using X-ray diffraction, extended X-ray absorption fine structure and 57Fe Mössbauer spectrometry, we have found a random disorder between the Fe and Ru sites. Contrary to theoretical predictions, the magnetic moment in this system is in fact contributed by Fe atoms, which calls into question the fundamental basis of the half-metallic character proposed for the FexRu2-xCrSi series. Our Mössbauer results also reveal a rare scenario in which key physical properties are intricately correlated with material chemistry in the form of random atomic disorder on a localised scale.
@@ I Introduction
Heusler alloys have recently garnered attention due to their diverse properties, including half-metallicity [10, 11, 12, 13, 14], spin gapless semiconductivity [15, 16, 17], spin-semimetals [18, 19], spin glass [20], re-entrant spin glass [21, 22, 23], bipolar magnetic semiconducting characteristics [24] and Curie temperature [25], etc. The exploration of new Heusler alloys is crucial for advancing this area of study, as despite theoretical predictions [26, 27, 28, 29] of numerous Heusler alloys possessing SGS features, only a limited number of experimental realizations have been achieved so far [30, 15, 16, 17]. Among them, Mn2CoAl [30], CoFeMnSi [15], CoFeCrGa [16], and CoFeMnSn [17] have been experimentally established as SGS materials, with the additional discovery of the fully compensated ferrimagnetic SGS CrVTiAl [31]. It is pertinent to note here that the most of the Heusler alloys often form with inherent and irrepressible structural disorder which is generally argued to be responsible for the nonconformity with the theoretically predicted HMF and SGS properties. Thus, the search for a compound with HMF and SGS properties, resistant to disorder, will be an effective solution for this type of situation.
@@
FexRu2-xCrSi is one such system where first principles band structure calculations predict robust HMF properties against structural disorder throughout the series [32, 33, 34]. The equiatomic member of the series, FeRuCrSi, has been additionally predicted to exhibit SGS characteristics [35, 36, 37]. The quaternary compound FeRuCrSi can be thought of as the intermediate between two ternary Heusler alloys Fe2CrSi [38] and Ru2CrSi [39]. Since Fe and Ru are isoelectronic atoms, the total valence electron count (VEC) of all members of FexRu2-xCrSi remain invariant (VEC = 26) irrespective to the value of $x$. Following the Slater-Pauling (S-P) rule [40, 41] for HMF Heusler alloy, the total magnetic moment for any member of the series is expected to be $M$ = (VEC-24) = 2 $\mu_{B}$/f.u. However, the Ru-rich end-composition, Ru2CrSi is experimentally found to undergo antiferromagnetic transition below $T_{N}\sim$14 K [39]. Subsequent theoretical calculations indeed show lower ground state energy for the antiferromagnetic spin arrangement in this compound [42]. As Fe concentration is gradually increased, the system initially transforms to a spin-glass state, before it finally settles to a ferromagnetic ground state for $x>0.3$ [43, 44, 45, 34, 46]. Fe2CrSi, the other-end member of the series, exhibits a Curie temperature at 520 K with saturation magnetic moment of 2 $\mu_{B}$/f.u., in line with the S-P rule [38]. Theoretical calculations suggest Cr-atoms are the major contributor of magnetism in this series of compounds, although this postulation is yet to be experimentally verified for compositions with $x\neq 0$. Interestingly, some of the experimental reports rather suggest that a linear variation of the magnetic moment occur with Fe substitution in FexRu2-xCrSi, contrary to the invariance of the magnetic moment expected from the S-P rule for HMF systems [43]. Additionally, even for FeRuCrSi, substantial variations in its basic characteristics like Curie temperature ($T_{\rm C}$) and saturation magnetization ($M_{sat}$) at low temperature have been reported in the literature [43, 36]. Thus, the members of FexRu2-xCrSi series not only sometimes fail to conform to the theoretical prediction, the experimental results on $T_{\rm C}$ and $M_{Sat}$ reported by different groups on the same compositions are also not reproducible. As mentioned earlier, since the Heusler alloys are known to form with inherent structural disorder, a priori one would expect that such variations in properties reported by various groups are related to different natures and extents of disorder present in the respective materials, despite having the same nominal composition. In this work, we have focussed on the equiatomic FeRuCrSi, and investigated the nature of disorders to explain the above-mentioned discrepancies among different experimental results as well as between experiment and theory. Very surprisingly, our investigation reveals that completely contrary to the theoretically established magnetic moment on Cr in FeRuCrSi, the magnetism is actually contributed by Fe atoms. The reported variations in $T_{\rm C}$ and $M_{sat}$ can also be explained by the local arrangement of Fe/Ru in each unit cells. Our work thus firmly establishes not only the role of anti-site disorders, but the effect of randomness of Fe/Ru distribution in this compound as well.
@@ II.1 Experimental
Extended X-ray absorption fine structure (EXAFS) measurements were conducted at the Energy Scanning EXAFS beamline (BL-9) at the Raja Ramanna Centre for Advanced Technology (RRCAT) in Indore, India, employing standard EXAFS measurement protocols [13, 21]. Magnetic measurements were carried out in the temperature range of 3–380 K under magnetic fields up to 70 kOe using a commercial SQUID-VSM instrument (Quantum Design Inc., USA). 57Fe transmission Mössbauer spectrometry was conducted at many different temperatures in the range of 405-77 K to investigate the local atomic environments of Fe atoms in the sample. An additional measurement under an externally applied magnetic field of 0.06 Tesla has also been carried out to investigate the influence of magnetic field in the Mössbauer spectra. The samples consisted of a thin powder layer containing about 5 mg Fe/cm2. Spectra were collected using an electromagnetic transducer with a triangular velocity profile and a 57Co source diffused into a Rh matrix within a bath cryostat for measurements below 300 K. The hyperfine structures were modelled using a least-square fitting procedure, involving quadrupolar doublets composed of Lorentzian lines, utilizing the in-house program ‘MOSFIT’. Isomer shift values were reported with reference to ${\alpha}$–Fe at 300 K, and the velocity was controlled using a standard of ${\alpha}$–Fe foil.
@@ II.2 Computational
The enthalpy of formation, electronic structure, and spin polarization at ground state were determined using Density Functional Theory (DFT) calculations with the projector augmented wave (PAW) method [48], as implemented in the Vienna ab initio simulation package (VASP) [49, 50]. The exchange-correlation was modelled by the generalized gradient approximation modified by Perdew, Burke, and Ernzerhof (GGA-PBE) [51]. Detailed information on calculation parameters can be found in previous articles [12, 13, 14, 21].
@@ III.1 Structure optimization and electronic structure calculations
Table 1: Calculated enthalpy of formation (${\Delta_{f}{H}}$) and magnetic moment for 3 different ordered atomic arrangements of FeRuCrSi and 2 different disordered structures. For these laters, the magnetic moment on $4c$ and $4$ corresponds to the average value of atoms in the respective ordered structure, e.g. Ru in $4c$ for Type-1.
sites distributions $\Delta_{f}{H}$ magnetic moment ($\mu_{B}$)
@@
Type-1 ordered Si Cr Ru Fe -33.63 0.00 1.98 -0.18 0.15 1.94
Type-1 disordered Si Cr (Ru,Fe) -34.13 0.00 1.90 -0.16 0.21 1.94
Type-2 ordered Si Ru Cr Fe -4.04 0.00 0.01 0.00 0.01 0.02
Type-3 disordered Si Fe Cr Ru -29.25 0.04 2.56 -1.18 0.30 1.73
Type-3 ordered Si Fe (Cr,Ru) -29.61 0.05 2.46 -0.87 0.19 1.85
A quaternary Heusler alloy, denoted as ${XX^{\prime}YZ}$, possesses four distinct crystallographic sites in space group $F\overline{4}3m$: 4a (0,0,0), 4b (0.5,0.5,0.5), 4c (0.25,0.25,0.25), and 4d (0.75,0.75,0.75). Generally, the sp-element Z occupies the 4a position, while the transition metals reside in the remaining three sites. Although there are six possible configurations, permutations of atoms in the 4c and 4d positions result in energetically equivalent configurations. Consequently, only the three distinct configurations (Type-1, Type-2, and Type-3) were considered, as shown in Fig.1 and Table.1. The relative stabilities of these configurations were assessed through DFT calculations by comparing their enthalpies of formation.
[…]
Consistent to that expected for HMF systems with VEC 26 [40, 52, 41], our calculation also yields the total magnetic moment to be 1.94 ${\mu_{B}}$. Moreover, atom-specific magnetic moments were computed and reported in Table. 1, and are in good agreement with earlier published reports [37, 35, 36]. The theoretical analysis suggests that Cr significantly contributes to the magnetic moment, and the moment on the Fe site rather exhibits an induced character. Similar behaviour has earlier been observed in FeMnVAl  [12] and FeMnVGa  [13] where magnetic atom Mn and nonmagnetic Fe are distributed randomly in the 4c and 4d sites, although in FeRuCrSi, Cr would occupy 4b sites whereas Ru and Fe would occupy 4c and 4d positions, respectively.
@@ III.2 X-ray diffraction
Figure 3: Rietveld refinement of the powder XRD pattern of FeRuCrSi (disordered structure) measured at room temperature. Inset (i) shows the magnified view of the XRD data for as-cast and annealed compound, where the impurity peaks in the annealed sample are marked by asterisk sign. Insets (ii) and (iii) show the magnified view of the Rietveld refinement results considering the ordered structure and disordered structure, respectively.
@@
We have carried out XRD measurements at room temperature for both as-cast as well as annealed sample. We found a few additional small peaks in the XRD patterns of the annealed sample which are absent in the XRD pattern of the as cast samples and also not conforming to the cubic character of Heusler alloy (inset (i) of Fig. 3). The result thus suggest that annealing tend to introduce secondary phase in the system. Consequently, all further measurements are carried out on the as-cast compound only. The Fig. 3 represent the Rietveld refinement of the XRD pattern of as-cast FeRuCrSi. Theoretical calculations suggests that ordered Type-1 structure possess minimum energy compared to other ordered structures. However, the earlier published experimental results on the crystal structure of FeRuCrSi are quite contradictory. Although some of the reports suggested the ordered LiMgPdSn structure-type [43], a few others also reported disordered atomic arrangements without determining the exact nature of disorder [36].
@@
Our attempt to fit the XRD pattern using the ordered Type-1 structure model however failed to properly address the superlattice (111) Bragg peak intensity (inset (ii) of Fig. 3). For Heusler systems, the (111) and (200) Bragg peak intensities are considered to be indicative of structural disorder present in the system. For example, in A2-type (random mixing of elements in 4a, 4b, 4c and 4d sites) disorder, both the peaks are found to be absent whereas in B2-type (cross-substitution between 4a & 4b and between 4c & 4d) only the (111) peak vanishes. In our case, both the 111 and 200 lines are present, thus ruling out A2 and B2-type disorder (inset (iii) of Fig. 3). Nonetheless, the intensity of 111 line is substantially reduced when compared to the theoretically estimated one for an ordered Type-1 structure. The reduction in intensity of the (111) peak, however, could be explained if the disorder is limited to the 4c and 4d sites only, where Fe and Ru atoms are distributed in 50:50 ratio (Type-1 disorder structure), averaging over the sample volume. However, it may be noted that the above structural model arising out of the XRD analysis, although is the simplest, yet certainly not unique. For example, a more complex structural model involving a combinations of Type-1 and Type-3 disorder with Si at 4a, Cr0.5Fe0.5 at 4b, Ru0.5Fe0.25Cr0.25 at 4c and Fe0.25Cr0.25Ru0.5 at 4d can also explain the XRD spectra with almost equal satisfaction. Nevertheless, as many other quaternary Heusler alloys, e.g. CoRuMnSi [53], FeMnVAl [12], FeMnVGa [13], NiRuMnSn [21] and FeRuVSi [54], etc. exhibiting similar XRD patterns have been established to form in the simpler disordered crystal structure by various theoretical and experimental tools, including neutron diffraction, we believe the same would also hold true for FeRuCrSi as well. The magnetic field dependent 57Fe Mösbauer spectroscopic data analysis of FeRuCrSi, presented later in Sec. III.5, indicates extremely local variation of hyperfine field which is more in consonance with the atomic disorder within the Type-I structure, rather than having a mixture of Type-I and Type-3 structure of much larger individual cluster sizes. The lattice parameter is found to be 5.757  Å, and matches with most of the other reported values [43, 37].
@@ III.3 EXAFS
The simultaneous fitting of Fe and Ru edges could only be achieved under the assumption of Fe/Ru mixing, confirming disorder in the examined compound based on the XRD data. The best-fit theoretical spectra and experimental data are presented in Fig. 5, and the corresponding parameters are listed in Table 2. The EXAFS estimates a ratio of Fe/Ru mixing equal to 59:41 and 45:55, by measuring from the Fe and Ru edges, respectively. The EXAFS data fittings align well with a disordered structure of Type-1, corroborating the XRD data analysis.
III.4 Magnetic properties
[…]